OpenStudy (anonymous):

The ratio of corresponding sides of similar triangles is four colon nine. If the longest side of the smaller triangle is thirty-six mm, what is the longest side of the larger triangle?

OpenStudy (anonymous):

The ratio of corresponding sides of similar triangles is 4:9. If the longest side of the smaller triangle is 36 millimeter, what is the longest side of the larger triangle?

OpenStudy (anonymous):

just figured i should have put it in # form...

OpenStudy (phi):

solve \[ \frac{4}{9}= \frac{36}{x} \]

OpenStudy (anonymous):

\[\frac{4}{9}=\frac{36}{x}\]

OpenStudy (anonymous):

thanks i couldnt get the right set up for this one....